OpenStudy (anonymous):

Simplify the following expression. 5square by 5 ÷ 5square by -3 a 5square by -15 b 5square by 2 c 5square by 8 d 5square by -8

OpenStudy (dangerousjesse):

I accidentally went back a page, sorry -.-

OpenStudy (dangerousjesse):

Simplify the following: \[5^2 \frac{5}{5^2} (-3)\]Express \(\frac{5}{5^2} (-3)\) as a single fraction. \[\frac{5}{5^2} (-3) = \frac{5 (-3)}{5^2}:\]\[5^2×\frac{5 (-3)}{5^2 }\]Express \(5^2×\frac{5 (-3)}{5^2}\) as a single fraction. \[5^2×\frac{5 (-3)}{5^2} = \frac{5 (-3) 5^2}{5^2}:\]\[\frac{5 (-3) 5^2}{5^2}\]Cancel common terms in the numerator and denominator of \(\frac{5 (-3) 5^2}{5^2}. \[\frac{5 (-3) 5^2}{5^2} = \frac{5^2}{5^2}×5 (-3) = 5 (-3):\]\[5 (-3) \]Multiply 5 and -3 together. \[5 (-3) = -15: \]

OpenStudy (dangerousjesse):

And it didn't even accept half of it. Dammit OP.

OpenStudy (dangerousjesse):

Cancel common terms in the numerator and denominator of \(\frac{5 (-3) 5^2}{5^2}.\) \[\frac{5 (-3) 5^2}{5^2} = \frac{5^2}{5^2}×5 (-3) = 5 (-3): \]\[5 (-3) \] Multiply 5 and -3 together. \[5 (-3) = -15: \]

OpenStudy (dangerousjesse):

Which leaves you with..?